OpenStudy (anonymous):

A 900kg safe is 1.6m above a heavy-duty spring when the rope holding the safe breaks. The safe hits the spring and compresses it 60cm..

OpenStudy (anonymous):

Answer I got -> 78400N/m confirm someone?

OpenStudy (anonymous):

... what is the question?

OpenStudy (anonymous):

are you asking for the spring constant k?

OpenStudy (anonymous):

knowing that the force that compresses the spring is the over-all wight W = mg of the safe... W = mg = 900kg(9.8m sec^-2) = 8820N = F = kx where x is the distance of compression = 60cm =0.6m then... k = F/x = 8820N/0.6m = 14700 N/m